OpenStudy (anonymous):

Mike wants to make meatloaf. His recipe uses a total of 3 pounds of meat. If he uses a 3 to 1 ratio of beef to pork, how much pork will he use? Enter your answer as a mixed number in simplest terms.

OpenStudy (nincompoop):

3:1 3 POUND OF BEEF:1 POUND OF PORK

OpenStudy (dan815):

3 pounds of meat in total is being used 3 to 1 ratio of beef means 3:1 or 3/(3+1) = 3/4 of the quantity is beed and 1/4th is pork

OpenStudy (mathstudent55):

A ratio of 3 to 1 beef to pork means that he uses 3 times as much beef as he uses pork.

OpenStudy (anonymous):

So my answer as a mixed number would be 1/4

OpenStudy (nincompoop):

SO BASED ON THAT, WE CAN SAY THAT IF THE TOTAL IS 3 LBS

OpenStudy (dan815):

1/4 * ( 3lb) is pork

OpenStudy (nincompoop):

DANG DAN DID IT ALREADY RESPONDING FROM PHONE SUCKS

TheSmartOne (thesmartone):

Nin... your caps lock button is pressed on =P

OpenStudy (dan815):

so 3/4 lb is pork

OpenStudy (nincompoop):

I THINK THE CONCEPT NEEDS TO BE ADDRESSED

OpenStudy (dan815):

do u understand O_O

OpenStudy (mathstudent55):

That means he will use 3x amount of beef, and 1x amount of pork. Notice that 3x and 1x are in the required 3:1 ratio of beef to pork. Now you need a total. 3x + 1x = 3 Solve for x Then 3x is the amount of beef he will use and x is the amount of pork he will use.

OpenStudy (dan815):